OpenStudy (anonymous):

Rafael has 1 black, 3 red, 7 green, 2 blue, and 2 yellow candies in his pocket. What is the probability he will pull a red one out of his pocket? the answer is 3 right??

OpenStudy (johnweldon1993):

Nooo not this one....15 total right? and there are 3 reds...so P(red) = 3 --- 15

OpenStudy (anonymous):

probabilities can never be greater than 1 probability of 1 means 100% surity of an event happening

OpenStudy (johnweldon1993):

and simplify that down a bit

OpenStudy (anonymous):

soo it is 1/3 ?!!?!?!?

OpenStudy (anonymous):

but i dont habe that option :(

OpenStudy (johnweldon1993):

No....not 5 -- 15 its 3 --- 15 simplify THAT

OpenStudy (anonymous):

have*

OpenStudy (anonymous):

2/15??

OpenStudy (johnweldon1993):

\[\frac{ 3 }{ 15 } = \frac{ 1 }{ 5 }\] 3 goes into 15 5 times right?

OpenStudy (anonymous):

right

OpenStudy (johnweldon1993):

Yeah...so 3/15 is simplified down to 1/5

OpenStudy (anonymous):

\[Probability = \frac{ number of favourable outcomes }{ Total number of outcomes }\] favpurable outcomes = no. of balls of one kind total no. of outcomes = total number of balls

OpenStudy (anonymous):

okaii